OpenStudy (anonymous):

1) A bungee jumper plummets from a high bridge to the river below and then bounces back over and over again. At time t seconds after her jump, her height H (in meters) above the river is given by the function:

OpenStudy (anonymous):

\[H(t) = 100+ 75e^{t/20}\cos((\pi/4)t) \]

OpenStudy (anonymous):

Find her height at the times indicated in the table.

OpenStudy (anonymous):

t: 0,1,2,4,6,8,12

OpenStudy (anonymous):

i just plug in right?

ganeshie8 (ganeshie8):

yes just plugin

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

np :) careful about cos ok cos(0) = 1 cos(pi/4), cos(pi/2), cos(pi).... compute them first

OpenStudy (anonymous):

cos(pi/4)=.53 aprox ?

ganeshie8 (ganeshie8):

cos(pi/4) = 1/sqrt(2)

OpenStudy (anonymous):

H(1) = 41.74?

OpenStudy (anonymous):

i just want to make sure to see if im doing it right

ganeshie8 (ganeshie8):

how did u get that, u should get more than 100 right

OpenStudy (anonymous):

i meant 141.74

ganeshie8 (ganeshie8):

\(H(t) = 100+ 75e^{t/20}\cos((\pi/4)t) \) put t= 1 \(H(1) = 100+ 75e^{1/20}\cos((\pi/4)1) \) \(H(1) = 100+ 75e^{1/20}\frac{1}{\sqrt{2}}\) http://www.wolframalpha.com/input/?i=100%2B+75e%5E%7B1%2F20%7D%5Cfrac%7B1%7D%7B%5Csqrt%7B2%7D%7D

ganeshie8 (ganeshie8):

H(1) = 155.75

ganeshie8 (ganeshie8):

did u get, H(0) = 175 ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good :) try the remaining values

OpenStudy (anonymous):

H(2)= 117.38

ganeshie8 (ganeshie8):

No. when u put t = 2, \(H(t) = 100+ 75e^{t/20}\cos((\pi/4)t) \) \(H(t) = 100+ 75e^{2/20}\cos((\pi/4)2) \) \(H(t) = 100+ 75e^{2/20}\cos(\pi/2) \)

ganeshie8 (ganeshie8):

wats cos (pi/2) ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so it would be 100

ganeshie8 (ganeshie8):

yes !

OpenStudy (anonymous):

and H(40 would be 8.5

OpenStudy (anonymous):

(4)**

ganeshie8 (ganeshie8):

\(H(t) = 100+ 75e^{t/20}\cos((\pi/4)t) \) put t=4 \(H(4) = 100+ 75e^{4/20}\cos((\pi/4)4) \) \(H(4) = 100+ 75e^{1/5}\cos(\pi) \) \(H(4) = 100+ 75e^{1/5}(-1) \) http://www.wolframalpha.com/input/?i=100%2B+75e%5E%7B1%2F5%7D%28-1%29

ganeshie8 (ganeshie8):

yes you're right lol 8.39 ~ 8.4

OpenStudy (anonymous):

lol yay but now im confused... H(8) is coming out to 211.75

ganeshie8 (ganeshie8):

\(H(8) = 100+ 75e^{8/20}\cos((\pi/4)8)\) \(H(8) = 100+ 75e^{8/20}\cos(2\pi)\) \(H(8) = 100+ 75e^{8/20} 1 \) http://www.wolframalpha.com/input/?i=100%2B+75e%5E%7B8%2F20%7D

ganeshie8 (ganeshie8):

wolfram says 211.89

OpenStudy (anonymous):

and for H(12) i got -36.5 is that even posssible considering bungee jumping?

ganeshie8 (ganeshie8):

lol he must be sinking in waters and dead be now :o let me think lol

OpenStudy (anonymous):

haha lol

ganeshie8 (ganeshie8):

okay wat that means is, our equation will not work for t >= 12 cuz, height -36.5 makes no sense. let me see if i cna quickly plot it

ganeshie8 (ganeshie8):

have a look at attached graph, after 10, the graph is becoming negative. so if the equation equation is really a practical one, the jumper wud die for sure in waters ;)

OpenStudy (anonymous):

hehe now we know more about bungee jumping :p

OpenStudy (anonymous):

quick question : y=-3sin pi(x) whats the amp and period? amp= -3 ...and what would the period be ? 2?

ganeshie8 (ganeshie8):

ahh yes be better cry mayday before t = 12 lol

ganeshie8 (ganeshie8):

*he

OpenStudy (anonymous):

haha :p

ganeshie8 (ganeshie8):

amp is distance from mean position. it can never be negative ok

ganeshie8 (ganeshie8):

y = -3sin(pix) amp = |-3| = +3

ganeshie8 (ganeshie8):

period is 2 correct !

OpenStudy (anonymous):

oh ok :) but how would i graph the period 2

ganeshie8 (ganeshie8):